a. Use complete square method to solve for x and sketch (x)=2x^2-9x+10

b. What is the range of the parabola?

c. What is the domain of the parabola?

d. Sketch your graph and label your vertex, x intercept and y intercept.

e. Use the Quadratic formula to verify the x intercepts.


1
Expert's answer
2021-04-29T12:36:35-0400

Solution.


"2x^2-9x+10=\newline= 2(x^2-2\cdot\frac{9}{4}x+\frac{81}{16})-\frac{81}{8}+10=\newline =2(x-\frac{9}{4})^2-\frac{1}{8}."

The graph of "y=2(x-\frac{9}{4})^2-\frac{1}{8}" is congruent to the basic parabola "y=x^2" , but is translated "\frac{9}{4}" units to the right, "\frac{1}{8}" units down and

compression 2 times to the y-axis.

Domain = R.

Range = "(-\frac{1}{8}, \infty)."

Vertex "(\frac{9}{4},-\frac{1}{8})."

Y intercept "(0,10)."

X intercept "2x^2-9x+10=0"

"D=81-4\u20222\u202210=1,"

"x_1=\frac{9-1}{4}=2,\newline
x_2=\frac{9+1}{4}=2.5."

So, "(2,0),(2.5,0)."